Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability-siblings

Expert replies
by lilu » Tue Mar 03, 2009 10:34 pm
In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?
5/21
3/7
4/7
5/7
16/21
Join the discussion
Source: — Problem Solving |

Re: Probability-siblings

by piyush_nitt » Wed Mar 04, 2009 3:16 am
lilu wrote:In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?
5/21
3/7
4/7
5/7
16/21
IMO E


There are 2 scenarios here :

1. Select 2 persons from group of 4 ,

Prob of selecting 2 person from group of 4 = 2C1 * 2C1 = 4

2. Select 1 from group of 4 and 1 from group of 3

Prob of selecting 1 person from group of 4 and 1 from group of 3 = 4C1*3C1 = 12

Total prob = 7C2 = 21

Therefore Ans is, Case 1 + Case 2 /21

16/21
Join the discussion

by venmic » Wed Apr 27, 2011 10:29 pm
can anyone give a better explanation for thsi please...
Join the discussion

by manpsingh87 » Wed Apr 27, 2011 11:18 pm
lilu wrote:In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?
5/21
3/7
4/7
5/7
16/21
probability of not selecting siblings= 1- (probability of selecting siblings);
we have 7 people 4 of which have exactly 1 sibling and 3 have exactly 2 siblings..!!!
let 7 people be p1,p2,p3,p4,p5,p6,p7; now out of these 7 peoples 4 people having exactly 1 sibling can be (p1&p2), (p3&p4); and people having exaclty two siblings can be (p5,p6,p7).

lets call them group 1(p1&2), group2(p3&p4) and group 3(p5,p6,p7)
out of 7 people any two people can be selected in 7C2 ways=21;

probability that two persons selected belongs to group1 =2C2=1;
probability that two persons selected belongs to group 2=2C2=1;
probability that two persons selected belongs to group 3=3C2=3;

hence the probability that selected persons will be sibling= 1+1+3/21=5/21;

required probability= 1-5/21=16/21; hence E
O Excellence... my search for you is on... you can be far.. but not beyond my reach!
Join the discussion

by Prashant Ranjan » Fri Jul 01, 2011 3:05 am
Consider the following sibling pairs S1, S2, and S3
S1 (A,B) S2(C,D) S3(E,F,G)

Now the probability of selecting A is 1C1. If A is selected then B cannot be selected and every other person can be selected. Thus the pairs can be AB, AC, AD, AE, AF, AG where the order of the selection doesn’t matter. So the probability of this selection is 1C1 * 5C1 / 7C2.
In the same way, B can also be paired with 5 other persons but not A.
BC, BD, BE, BF, BG. So the selection probability is 1C1*5C1/7C2.
Now consider the selection of C. CE, CF, CG i.e. 1C1*3C1/7C2 (since CA and CB have already been counted).
Same with D. DE, DF, DG i.e. 1C1*3C1/7C2 (Since DA, DB have already been counted).
However when we select E we cannot select the persons F and G, since they are kins. Also we cannot select any other person from the 4 people since that pair has already been counted above.
Same goes for F and G. Thus the total probability of selection is
5C1/7C2 + 5C1/7C2 + 3C1/7C2 + 3C1/7C2 = 16C1/7C2 = 16/21
Join the discussion